a rectangle has an area of 12/36 sq unit. provide the dimensions of three rectangles that would have this area
Answer by Alan3354(69443)

a rectangle has an area of 12/36 sq unit. provide the dimensions of three rectangles that would have this area
=============
Area = L*W = 1/3
----
L = 1, W = 1/3
L = pi, W = (1/3)/pi
L = 1000, W = 1/3000